There are 200 questions on a 3 hr examination.Among these questions are 50 mathematics problems.It is suggested that twice as much time be spent on each maths problem as for each other question.How many minutes should be spent on mathematics problems
Answer by nerdybill(7384)

There are 200 questions on a 3 hr examination.Among these questions are 50 mathematics problems.It is suggested that twice as much time be spent on each maths problem as for each other question.How many minutes should be spent on mathematics problems
.
First, since we're talking about minutes, convert 3hr into minutes:
3*60 = 180 minutes
.
Let x = minutes spent on a non-math problem
then
2x = minutes spent on a math problem
.
x(200-50) + 2x(50) = 180
x(150) + 2x(50) = 180
150x + 100x = 180
250x = 180
x = 0.72 minutes
.
Time to spend on math problems:
2x(50)=100x = 100(0.72) = 72 minutes